Magnetic fields at the surface of the sun can be s
Answer #1

magnetic field energy density is given by

energy 1 B2 ,-volume 2 μ По-

total volume of mass released=mass/density=3.333*10^18 m^3

total energy=0.5 *B^2*volume/uo=0.5*m*v^2

solving this we get B=11.64 T